OpenStudy (anonymous):

find the laplace transform of y''+y=sin(t) y(0)=y'(0)=0

OpenStudy (usukidoll):

 use Laplace and plug those conditions in

OpenStudy (usukidoll):

and I have yet to learn Laplace...getting there

OpenStudy (unklerhaukus):

\[\qquad y''+y=\sin (t)\] \[\quad\mathcal L\big\{y''+y\big\}=\mathcal L\big\{\sin (t) \big\}\] \[\mathcal L\big\{y''\}+\mathcal L\big\{y\big\}=\mathcal L\big\{\sin (t) \big\}\]

OpenStudy (unklerhaukus):

\[\mathcal L\big\{y''\}=s^2Y(s)-sy(0)-y'(0)\] \[\mathcal L\big\{y\big\}=Y(s)\] \[\mathcal L\big\{\sin (nt) \big\}=\frac{n}{s^2+n^2}\]

OpenStudy (unklerhaukus):

\[\mathcal L\big\{y''\}+\mathcal L\big\{y\big\}=\mathcal L\big\{\sin (t) \big\}\\ [s^2Y(s)−sy(0)−y'(0)]+[Y(s)]=\frac{n}{s^2+n^2}\\ (s^2+1)Y(s)=\frac{n}{s^2+n^2}\\ Y(s)=\frac{n}{(s^2+n^2)}\times\frac1{(s^2+1)}\\ y(t)=\mathcal L^{-1}\Big\{\frac{n}{(s^2+n^2)}\times\frac1{(s^2+1)}\Big\}\]